### 10 Jan 2024 v1

# A proof of Audenaert-Kittaneh’s Conjecture

Abstract
We show that Audenaert-Kittaneh’s Conjecture related to $p$-Schatten class.
keywords:
Schatten $p$-norm, inequality
1 Introduction
The classical Clarkson’s inequalities C36 for the Lebesgue spaces $L_{p}$, and their non-commutative analogues for the Schatten trace ideals, play an important role in analysis, operator theory, and mathemetical physics. They have been generalised in various directions. Among these versions for more-general symmetric norms BH88 , for operators by $n$-tuples BK04 ; HK08 and for the Haagerup $L_{p}$ spaces FK86 , as well as refinements BCL94 . Let $\mathbb{B}(\mathscr{H})$ denote the algebra of all bounded linear operators acting on a complex separable Hilbert space $\mathscr{H}$. If $X\in\mathbb{B}(\mathscr{H})$ is compact, we denote by $\{s_{j}(X)\}$ the sequence of decreasingly ordered singular values of $X$. For $p>0$, let
$\|X\|_{p}=\left[(s_{j}(X))^{p}\right]^{1/p}=\left(\text{tr }\left|X\right|^{p}\right)^{\frac{1}{p}},$
[…]
$\mathbb{B}_{p}(\mathscr{H})=\{X\in\mathbb{B}(\mathscr{H}):\|X\|_{p}<\infty\},$
which is called the $p$-Schatten class of $\mathbb{B}(\mathscr{H})$; cf. GK69 .
Clarkson’s inequalities for operators in $\mathbb{B}_{p}(\mathscr{H})$ (see Mc67 ) assert that for $0<p\leq 2$
$2^{p-1}(\|A\|_{p}^{p}+\|B\|_{p}^{p})\leq\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}\leq 2(\|A\|_{p}^{p}+\|B\|_{p}^{p}),$ (1)
and for $p\geq 2$
$2(\|A\|_{p}^{p}+\|B\|_{p}^{p})\leq\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}\leq 2^{p-1}(\|A\|_{p}^{p}+\|B\|_{p}^{p}).$ (2)
For $p=2$ both inequalities (1) and 2 reduce to the parallelogram law
$\|A-B\|_{2}^{2}+\|A+B\|_{2}^{2}=2(\|A\|_{2}^{2}+\|B\|_{2}^{2}).$
On the other hand, McCarthy Mc67 obtained for $1<p\leq 2$ the following inequality
$2(\|A\|_{p}^{p}+\|B\|_{p}^{p})^{\frac{q}{p}}\leq\|A+B\|_{p}^{q}+\|A-B\|_{p}^{q}$ (3)
and for $p\geq 2$
$\|A+B\|_{p}^{q}+\|A-B\|_{p}^{q}\leq 2(\|A\|_{p}^{p}+\|B\|_{p}^{p})^{\frac{q}{p}}$ (4)
where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
One natural generalizations of (1) and (2) are as follows:
Theorem 1.1.
(HK08, , Hirazallah & Kittaneh) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $0<p\leq 2$
$n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq\sum_{i=1}^{n}\|A_{i}\|^{p}_{p},$ (5)
and for $p\geq 2$
$n\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}.$ (6)
In view of the inequalities (3) and (4), it is reasonable to make the following conjecture concerning generalizations (along the lines of the inequalities (5) and (6) of the inequalities (1) and (2)) to $n$-tuples of operators.
Conjecture 1.2.
(AK12, , Audenaert & Kittaneh) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\right)^{\frac{q}{p}}$
and for $p\geq 2$
$n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$
where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
In CM16 , Conde and Moslehian obtained a weaker version of Conjecture 1.2 with a weaker coefficient $n^{\frac{q}{2}}$ instead of $n$, the conjecture 1.2 remains open.
Theorem 1.3.
(CM16, , Theorem 3.2) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then
$n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$
for $p\in[2,\infty);\frac{1}{p}+\frac{1}{q}=1$,
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}$
for $p\in(1,2];\frac{1}{p}+\frac{1}{q}=1$.
In this paper, we give a proof of Audenaert-Kittaneh’s Conjecture along a completely different path from Conde and Moslehian.
2 Proof of Audenaert-Kittaneh’s Conjecture
In this section, we give a proof of Audenaert-Kittaneh’s Conjecture 1.2, the ideas are the same as those used in FK86 ; BK04 .
Lemma 2.4 (Hadamard three-lines theorem, see U08 ).
Let $f(z)$ be a bounded function of $z=x+iy$ defined on the strip
$\{x+iy:a\leq x\leq b\},$
holomorphic in the interior of the strip and continuous on the whole strip. If
$M(x):=\sup\limits_{y}\left|f(x+iy)\right|,$
then $\log M(x)$ is a convex function on $[a,b]$. In other words, if $x=ta+(1-t)b$ with $0\leq t\leq 1$, then
$M(x)\leq M(a)^{t}M(b)^{1-t}.$
Lemma 2.5.
Let $A_{1},\ldots,A_{n}\in B_{p}(\mathscr{H})$ for some $1<p\leq 2$. Denote $B=\sum\limits_{k=1}^{n}A_{k},B_{i,j}=A_{i}-A_{j},1\leq i<j\leq n$. Let $Y,Y_{i,j}$ be operators in the dual class $B_{q}(\mathscr{H})$. Then
|tr ($\left|\text{tr }\left(YB+\sum_{1\leq i<j\leq n}Y_{i,j}B_{i,j}\right)\right|\leq n^{\frac{1}{q}}\left(\sum_{k=1}^{n}\|A_{k}\|_{p}^{p}\right)^{\frac{1}{p}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{p}}.$ (7)
Proof.
Let $A_{k}=\left|A_{k}\right|W_{k},Y=V\left|Y\right|$ and $Y_{i,j}=V_{i,j}\left|Y_{i,j}\right|$ be right, left and left polar decompositions of $A_{k},Y$ and $Y_{i,j}$, respectively. Here, $W_{j},V$ and $V_{i,j}$ are partial isometries. We have $11/2\leq 1/p<$ 1. For the complex variable $z=x+iy$ with $1/2\leq x\leq 1$, let
$A_{k}(z)$ $=$ $\left|A_{k}\right|^{pz}W_{k};$
$B(z)$ $=$ $\sum\limits_{k=1}^{n}A_{k}(z);$
$B_{i,j}(z)$ $=$ $A_{i}(z)-A_{j}(z);$
$Y(z)$ $=$ $\|Y\|_{q}^{pz-q(1-z)}V\left|Y\right|^{q(1-z)};$
$Y_{i,j}(z)$ $=$ $\|Y_{i,j}\|_{q}^{pz-q(1-z)}V_{i,j}\left|Y_{i,j}\right|^{q(1-z)}.$
Note that $A_{k}(1/p)=A_{k},Y(1/p)=Y$ and $Y_{i,j}(1/p)=Y_{i,j}$. Let
$f(z)=\text{tr }\left(Y(z)B(z)+\sum_{1\leq i<j<n}Y_{i,j}(z)B_{i,j}(z)\right).$
The left-hand side of (7) is $\left|f(1/p)\right|$. We can estimate this if we have bounds for $\left|f(z)\right|$ at $x=1/2$ and $x=1$. If $x=1$, we have
|tr $\left|\text{tr }Y(z)A_{k}(z)\right|$ $=$ $\|Y\|_{q}^{p}\left|\text{tr }V\left|Y\right|^{-iqy}\left|A_{k}\right|^{p(1+iy)}W_{k}\right|;$
|tr $\left|\text{tr }Y_{i,j}(z)A_{k}(z)\right|$ $=$ $\|Y_{i,j}\|_{q}^{p}\left|\text{tr }V_{i,j}\left|Y_{i,j}\right|^{-iqy}\left|A_{k}\right|^{p(1+iy)}W_{k}\right|.$
Using the information that for any operator $T$,
|tr $\left|\text{tr }T\right|\leq\|T\|_{1}\text{ and }\|XTY\|\leq\|X\|_{\infty}\|T\|\|X\|_{\infty}$
[…]
|tr $\left|\text{tr }Y(z)A_{k}(z)\right|$$\leq$ $\|Y\|_{q}^{p}\left\|A_{k}\right\|_{p}^{p};$
|tr $\left|\text{tr }Y_{i,j}(z)A_{k}(z)\right|$$\leq$ $\|Y_{i,j}\|_{q}^{p}\left\|A_{k}\right\|_{p}^{p}.$
Hence
$\left|f(z)\right|$ $=$ |tr ($\left|\text{tr }\left(Y(z)B(z)+\sum_{1\leq i<j\leq n}Y_{i,j}(z)B_{i,j}(z)\right)\right|$ (8)
$=$ |tr ($\left|\text{tr }\left(Y(z)\sum_{k=1}^{n}A_{k}(z)+\sum_{1\leq i<j\leq n}Y_{i,j}(z)\left(A_{i}(z)-A_{j}(z)\right)\right)\right|$
$=$ |tr $\left|\text{tr }\sum_{k=1}^{n}\left(Y(z)-\sum\limits_{i=1}^{k-1}Y_{i,k}(z)+\sum\limits_{i=k+1}^{n}Y_{k,i}(z)\right)A_{k}(z)\right|$
$\leq$ $\sum_{k=1}^{n}\left(\left|\text{tr }Y(z)A_{k}(z)\right|+\sum\limits_{i=1}^{k-1}\left|\text{tr }Y_{i,k}(z)A_{k}(z)\right|+\sum\limits_{i=k+1}^{n}\left|\text{tr }Y_{k,i}(z)A_{k}(z)\right|\right)$
$\leq$ $\sum_{k=1}^{n}\left(\|Y\|_{q}^{p}+\sum\limits_{i=1}^{k-1}\|Y_{i,k}\|_{q}^{p}+\sum\limits_{i=k+1}^{n}\|Y_{k,i}\|_{q}^{p}\right)\|A_{k}\|_{p}^{p}$
$\leq$ $\left(\left\|Y\right\|_{q}^{p}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{q}^{p}\right)\left(\sum_{k=1}^{n}\|A_{k}\|_{p}^{p}\right)$
when $x=1$. When $x=1/2$, the operators $A_{k}(z)$ and $Y_{i,j}(z)$ are in $\mathbb{B}_{2}(\mathscr{H})$ and
$\left|f(z)\right|$$\leq$ |tr $\left|\text{tr }Y(z)B(z)\right|+\sum_{1\leq i<j\leq n}\left|\text{tr }Y_{i,j}(z)B_{i,j}(z)\right|$
$\leq$ $\|Y(z)\|_{2}\|B(z)\|_{2}+\sum_{1\leq i<j\leq n}\|Y_{i,j}(z)\|_{2}\|B_{i,j}(z)\|_{2}$
$\leq$ $\left(\left\|Y(z)\right\|_{2}^{2}+\sum_{1\leq i<j\leq n}\|Y_{i,j}(z)\|_{2}^{2}\right)^{\frac{1}{2}}\left(\left\|B(z)\right\|_{2}^{2}+\sum_{1\leq i<j\leq n}\|B_{i,j}(z)\|_{2}^{2}\right)^{\frac{1}{2}}$
$=$ $n^{\frac{1}{2}}\left(\left\|Y(z)\right\|_{2}^{2}+\sum_{1\leq i<j\leq n}\|Y_{i,j}(z)\|_{2}^{2}\right)^{\frac{1}{2}}\left(\sum_{i=1}^{n}\|A_{i}(z)\|_{2}^{2}\right)^{\frac{1}{2}}.$
The equality at the last step is a consequence of Theorem 1.3, specialised to the case $p=2$. Note that when $x=1/2$, we have $\|A_{k}(z)\|_{2}^{2}=\|A_{k}\|_{p}^{p}$, and $\|Y_{i,j}(z)\|_{2}^{2}=\|Y_{i,j}\|_{q}^{p}$. Hence
$\left|f(z)\right|\leq n^{\frac{1}{2}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{2}}\left(\sum_{i=1}^{n}\|A_{i}(z)\|_{p}^{p}\right)^{\frac{1}{2}},$ (9)
when $x=1/2$. Let $M_{1},M_{2}$ be the right sides of (8), (9), repectively. Then by Lemma 2.4, we have, for $11/2\leq 1/p<$ 1,
$\left|f(1/p)\right|\leq M_{1}^{2(1/p-1/2)}M_{2}^{2(1-1/p)}.$
This gives (7). ∎
Proof of Audenaert-Kittaneh’s Conjecture: Now, let $B=U\left|B\right|,B_{i,j}=U_{i,j}\left|B_{i,j}\right|$ be a polar decomposition, and let
$Y=\left\|B\right\|_{p}^{q-p}\left|B\right|^{p-1}U^{*},Y_{i,j}=\left\|B_{i,j}\right\|_{p}^{q-p}\left|B_{i,j}\right|^{p-1}U_{i,j}^{*}.$
It is easy to see that
$\text{tr }YB=\left\|B\right\|_{p}^{q}=\left\|Y\right\|_{q}^{p},\text{tr }Y_{i,j}B_{i,j}=\left\|B_{i,j}\right\|_{p}^{q}=\left\|Y_{i,j}\right\|_{q}^{p}.$
So from Lemma 2.5, we get
$\|B\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|B_{i,j}\|_{p}^{q}\leq n^{\frac{1}{q}}\left(\sum_{i=1}^{n}\left\|A_{i}\right\|_{p}^{p}\right)^{1/p}\left(\|B\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|B_{i,j}\|_{p}^{q}\right)^{1/p}.$
This is the same as saying that
$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}},1<p\leq 2.$
This proves the conjecture for $1<p\leq 2$. The reverse inequality for $2\leq p<\infty$ can be obtained from this by a duality argument.
Acknowledgments
I heartily thank my advisor Minghua Lin for his guidance in my study and care in my life.

### 26 Feb 2024 v2

@@
-A proof of Audenaert-Kittaneh’s Conjecture
+
 Abstract
-We show that Audenaert-Kittaneh’s Conjecture related to $p$-Schatten class.
+By using Hadamard’s 3-lines theorem for a certain analytic function defined in terms of the trace, we show that Audenaert-Kittaneh’s conjecture related to $p$-Schatten class.
 keywords:
-Schatten $p$-norm, inequality
+Schatten $p$-norm, inequality, Hadamard’s 3-lines theorem, trace function
+journal: Journal of Functional Analysis
 1 Introduction
-The classical Clarkson’s inequalities C36 for the Lebesgue spaces $L_{p}$, and their non-commutative analogues for the Schatten trace ideals, play an important role in analysis, operator theory, and mathemetical physics.
+The classical Clarkson’s inequalities C36 for the Lebesgue spaces $L_{p}$, and their non-commutative analogues for the Schatten trace ideals, play an important role in analysis, operator theory, and mathematical physics.
 They have been generalised in various directions.
@@
 $2(\|A\|_{p}^{p}+\|B\|_{p}^{p})\leq\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}\leq 2^{p-1}(\|A\|_{p}^{p}+\|B\|_{p}^{p}).$ (2)
-For $p=2$ both inequalities (1) and 2 reduce to the parallelogram law
+For $p=2$ both inequalities (1) and (2) reduce to the parallelogram law
 $\|A-B\|_{2}^{2}+\|A+B\|_{2}^{2}=2(\|A\|_{2}^{2}+\|B\|_{2}^{2}).$
+In BCL94 , Ball, Carlen and Lieb explained the following theorem fully.
+Theorem 1.1 (Optimal 2-uniform convexity).
+For $1\leq p\leq 2$, the inequality
+$\left(\dfrac{\|A+B\|_{p}^{p}+\|A-B\|_{p}^{p}}{2}\right)^{\frac{2}{p}}\geq\|A\|_{p}^{2}+(p-1)\|B\|_{p}^{2}$ (3)
+holds in the following case:
+  1. 1.
+    $A$ and $B$ are functions in $L_{p}$.
+  2. 2.
+    $A$ and $B$ are operators in $B_{p}(\mathscr{H})$.
+If $2\leq p\leq\infty$, the inequality is reversed.
+We note that validity of the inequality (3) implies 2-uniform convexity of $L_{p}$ and $C_{p}$.
+When $1\leq p\leq 2$, the function $f(t)=t^{\frac{p}{2}}$ is concave.
+So we can get the $\frac{p}{2}$-th power of the right side of (3)
+$(\|A\|_{p}^{2}+(p-1)\|B\|_{p}^{2})^{\frac{p}{2}}$$\geq$ $2^{\frac{p}{2}-1}\|A\|_{p}^{p}+2^{\frac{p}{2}-1}(p-1)^{\frac{p}{2}}\|B\|_{p}^{p}$
+$\geq$ $2^{\frac{p}{2}-1}(\|A\|_{p}^{p}+\|B\|_{p}^{p}).$
+It implies that (3) is sharper than the first inequality of (1).
 On the other hand, McCarthy Mc67 obtained for $1<p\leq 2$ the following inequality
+$\|A+B\|_{p}^{q}+\|A-B\|_{p}^{q}\leq 2(\|A\|_{p}^{p}+\|B\|_{p}^{p})^{\frac{q}{p}}$ (4)
+and for $p\geq 2$
 $2(\|A\|_{p}^{p}+\|B\|_{p}^{p})^{\frac{q}{p}}\leq\|A+B\|_{p}^{q}+\|A-B\|_{p}^{q}$ (5)
-and for $p\geq 2$
-$\|A+B\|_{p}^{q}+\|A-B\|_{p}^{q}\leq 2(\|A\|_{p}^{p}+\|B\|_{p}^{p})^{\frac{q}{p}}$ (4)
 where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
-One natural generalizations of (1) and (2) are as follows:
+One natural generalizations of partial (1) and (2) are as follows:
 Theorem 1.2.
 (HK08, , Hirazallah & Kittaneh) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $0<p\leq 2$
-$n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq\sum_{i=1}^{n}\|A_{i}\|^{p}_{p},$ (5)
+$n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p},$ (6)
 and for $p\geq 2$
-$n\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\leq\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}.$ (6)
-In view of the inequalities (3) and (4), it is reasonable to make the following conjecture concerning generalizations (along the lines of the inequalities (5) and (6) of the inequalities (1) and (2)) to $n$-tuples of operators.
-Conjecture 1.2.
-(AK12, , Audenaert & Kittaneh) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
-$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{q}_{p}\right)^{\frac{q}{p}}$
+$\|\sum_{i=1}^{n}A_{i}\|_{p}^{p}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{p}\leq n^{p-1}\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}.$ (7)
+In view of the inequalities (4) and (5), it is reasonable to corresponding generalizations (along the lines of the inequalities (6) and (7) of the partial inequalities (1) and (2)) to $n$-tuples of operators.
+More precisely, in (AK12, , Section 8.1) entitled “Clarkson inequalities for several operators”, the authors presented the following conjecture:
+Audenaert-Kittaneh’s Conjecture.
+Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
+$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}},$ (8)
 and for $p\geq 2$
-$n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$
+$n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$ (9)
 where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
-In CM16 , Conde and Moslehian obtained a weaker version of Conjecture 1.2 with a weaker coefficient $n^{\frac{q}{2}}$ instead of $n$, the conjecture 1.2 remains open.
+The above inequalities Audenaert and Kittaneh conjectured is a crucial piece of the Clarkson-type inequality puzzle.
+Notice that both of (8) and (9) become equalities for $A_{1}=\ldots=A_{n}$, so these two inequalities are sharp.
+But little progress has been made in the decade since the conjecture was first proposed.
+It is worth mentioning that Conde and Moslehian obtained a weaker version of Audenaert-Kittaneh’s conjecture Audenaert-Kittaneh’s Conjecture with a weaker coefficient $n^{\frac{q}{2}}$ instead of $n$ in 2016.
 Theorem 1.3.
-(CM16, , Theorem 3.2) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then
+(CM16, , Theorem 3.2) Let $A_{1},\ldots,A_{n}\in\mathbb{B}_{p}(\mathscr{H})$. Then for $1<p\leq 2$
+$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}},$
+and for $p\geq 2$
 $n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}\leq\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}$
-for $p\in[2,\infty);\frac{1}{p}+\frac{1}{q}=1$,
-$\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n^{\frac{q}{2}}\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}}$
-for $p\in(1,2];\frac{1}{p}+\frac{1}{q}=1$.
+where $p,q>0$ and $\frac{1}{p}+\frac{1}{q}=1$.
 In this paper, we give a proof of Audenaert-Kittaneh’s Conjecture along a completely different path from Conde and Moslehian.
 2 Proof of Audenaert-Kittaneh’s Conjecture
-In this section, we give a proof of Audenaert-Kittaneh’s Conjecture 1.2, the ideas are the same as those used in FK86 ; BK04 .
+In this section, we give a proof of Audenaert-Kittaneh’s conjecture Audenaert-Kittaneh’s Conjecture.
+First, we introduced Hadamard three-lines theorem, which appears in the classic complex analysis textbook.
 Lemma 2.4 (Hadamard three-lines theorem, see U08 ).
@@
 $M(x)\leq M(a)^{t}M(b)^{1-t}.$
+Next, we give an elegant use of Hadamard’s three-lines lemma 2.4 for a certain analytic function defined in terms of the trace.
 Lemma 2.5.
@@
 Hence
-$\left|f(z)\right|\leq n^{\frac{1}{2}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{2}}\left(\sum_{i=1}^{n}\|A_{i}(z)\|_{p}^{p}\right)^{\frac{1}{2}},$ (9)
+$\left|f(z)\right|\leq n^{\frac{1}{2}}\left(\left\|Y\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|Y_{i,j}\|_{p}^{q}\right)^{\frac{1}{2}}\left(\sum_{k=1}^{n}\|A_{k}\|_{p}^{p}\right)^{\frac{1}{2}},$ (12)
 when $x=1/2$. Let $M_{1},M_{2}$ be the right sides of (11), (12), repectively. Then by Lemma 2.4, we have, for $11/2\leq 1/p<$ 1,
@@
 $\left\|\sum_{i=1}^{n}A_{i}\right\|_{p}^{q}+\sum_{1\leq i<j\leq n}\|A_{i}-A_{j}\|_{p}^{q}\leq n\left(\sum_{i=1}^{n}\|A_{i}\|^{p}_{p}\right)^{\frac{q}{p}},1<p\leq 2.$
-This proves the conjecture for $1<p\leq 2$. The reverse inequality for $2\leq p<\infty$ can be obtained from this by a duality argument.
+This proves the conjecture for $1<p\leq 2$. The reverse inequality for $2\leq p<\infty$ can be obtained from this by a duality argument. ∎
+Conflict of interest and data availability statement
+The author reports no conflicting interests or funding. Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
 Acknowledgments
-I heartily thank my advisor Minghua Lin for his guidance in my study and care in my life.
+The author heartily thanks professor Eric A.
+Carlen for several valuable suggestions and comments, including introducing Hanner’s inequality and pointing out the result in this paper is best possible etc.
+The author would like to his advisor Minghua Lin for guidance in his study and care in his life.